OpenStudy (anonymous):

Cooks use baking soda to make cakes light and fluffy. You might have used baking soda yourself. Baking soda is NaHCO3. A 0.1 M solution of baking soda and water has a [H+] of about 4.0 X 10^-9. (You may prefer to think of the Hydronium ion concentration, [H3O+], as 4.0 X 10-9.) write the formula for the calculation of pH, and then show each job as you calculate the pH of 0.1 M solution of baking soda.

OpenStudy (micahm):

www.chemistryland.com/CHM130S/LabHelp/Experiment12/Exp12.html

OpenStudy (cuanchi):

NaHCO3 -> Na+ + HCO3- (dissociate completely it is a ionic compound) HCO3- + H20 -> H2CO3 + OH- Kb = Kw/Ka Ka= 4.3 x 10-7 (carbonic acid Ka) (Hydrolysis ) calculate [OH-] whit the value of Kb and a ICE table and the equilibrium expression or by using the approximation formula for the x is too small \[[OH ^{-}]=\sqrt{\frac{ Kw }{ Ka }\times Cs}\] pOH= -log [OH-] pH= 14-pOH